If I want a fund of $30,000 in 14 years earning 5% quarterly interest... how much do I need to deposit now? How would I go about setting up this equation?
Answer by richwmiller(17219)


A=p(1+i)^n
i=.05/4=.0125
n=14*4=56
A=30,000
30000=p(1+(.00125))^56=27973